OpenStudy (anonymous):

Can someone help me graph this? : (x^2/25)+((y-2)^2/9)=1 I want to know the center, vertices, co vertices, and foci. I thought I had it right but checked on WolfrumAlpha and my graph didn't look the same. Help? :)

terenzreignz (terenzreignz):

General equation of an ellipse :) \[\Large \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]Now kindly identify the values of h,k,a, and b ^.^

OpenStudy (anonymous):

h: 0 k: 2 a: 5 b: 3

terenzreignz (terenzreignz):

That was fast :) Now, the centre is to be found at the point (h,k) So that clears ONE thing up ^.^

OpenStudy (anonymous):

ohhhhh... nevermind!!! I got it! Just forgot the general equation! :)

OpenStudy (anonymous):

Thank you!

terenzreignz (terenzreignz):

Glad that worked out :)